Give an example of two metric on R2 which are equivalent ,substantiate your claim


1
Expert's answer
2021-07-23T12:03:27-0400

Solution:

Consider two metric spaces on R2;

1.Euclidean metric de

"d_e(\underline{x},\underline{y})=\sqrt{(x_1-y_1)^2+(x_2-y_2)^2}"

2. The max-metric dm

"d_m(\underline{x},\underline{y})="max{ |x1-y1|,|x2-y2|}

We have;

"d_m(\underline{x},\underline{y})"=max{|x1-y1|,|x2-y2|}"\leq" "\sqrt{(x_1-y_1)^2+(x_2-y_2)^2}" ="d_e(\underline{x},\underline{y})"

Also;

"d_e(\underline{x},\underline{y})=\sqrt{(x_1-y_1)^2+(x_2-y_2)^2}=\sqrt{(|x_1-y_1|)^2+(|x_2-y_2|)^2}"

"d_e(\underline{x},\underline{y})\leq\sqrt{d_m^2(\underline{x},\underline{y})+d_m^2(\underline{x},\underline{y})}" ="\sqrt{2}d_m(\underline{x},\underline{y})"

Hence,

"d_m(\underline{x},\underline{y})\leq d_e(\underline{x},\underline{y})\leq \sqrt{2}d_m(\underline{x},\underline{y})"

for all x,y"\epsilon" X.

Hence, metrics dm and de are equivalent.